is this endo or exo?Use the following bond energies to decide if the reaction of CO2 +2H2 --> CH4 + O2 is endo or exo thermic. C--H 414 kJ/mol S
Answer #1

Co2 + 2H2 - CH 4 to o n DH = ? o=cro + 2 H -H -Ć -H + 0 =0 7 DH = bond enery of reactents- products = 2 x Bof- of Cro + 2x Br